OpenStudy (anonymous):

Which of the following is the equation of a line with a slope of 2 that passes through (1, -3)?

OpenStudy (anonymous):

This is as far as I got: y - y1 = m(x - x1) y+3=2(x-1)

OpenStudy (mathmale):

ILMC: Your equation is perfect. You could, if you wish, solve it for y. Either your equation, as is, or after you solve it for y, should match one of the answer options (which you haven't listed, by the way).

OpenStudy (anonymous):

Thanks :) Yeah here are the options: a. y = -2x+5 b. y = 2x-5 c. y = 2x+1 d. y = 2x+3

OpenStudy (mathmale):

ILMC: start with the equation you already have. subtract 3 from both sides. (Why?)

OpenStudy (anonymous):

So, y=-1(x-1)?

OpenStudy (mathmale):

Thanks for trying this. There are rules in algebra called "order of operations." We must multiply and divide before adding and subtracting. So, starting with your y+3=2(x-1), please distribute that multiplier of 2 on the right side. In other words, multiply what's inside the parentheses by 2 BEFORE you subtract 3 from both sides of the equation.

OpenStudy (anonymous):

So y=2x-5?

OpenStudy (mathmale):

Perfect. Keep in mind those "order of operations" rules. Please do an internet search for "order of operations" and see what shows up. Otherwise, your work is perfect.

OpenStudy (anonymous):

Thank you c:

OpenStudy (mathmale):

My great pleasure. And thanks for the medal!